### v1

@@ A.1.2 Exploiting sparse matrix operations for updating variational parameters
$\Lambda_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)~{}$ $=~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left[g_{ij}\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right\}+g_{ji}\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right\}\right]$
$\hskip 14.22636pt+\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left[(1-g_{ij})\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right\}+(1-g_{ji})\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right\}\right]$
$=\sum_{\bm{x}\in\mathcal{X}}\sum_{j\neq i}^{N}\sum_{l=1}^{K}g_{ij}\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x})}{\pi_{kl}^{(s)}(0,\bm{0})}\right\}$
$\hskip 14.22636pt+g_{ji}\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x})}{\pi_{lk}^{(s)}(0,\bm{0})}\right\}+(1-g_{ij})\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\,\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x})}{\pi_{kl}^{(s)}(0,\bm{0})}\right\}$
$\hskip 14.22636pt+(1-g_{ji})\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\,\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x})}{\pi_{lk}^{(s)}(0,\bm{0})}\right\}$
@@
$\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\xi}^{(s)}\right)\coloneqq\begin{bmatrix}\log\frac{\pi_{11}(d,\bm{x})}{\pi_{11}(d,\bm{0})}&\log\frac{\pi_{12}(d,\bm{x})}{\pi_{12}(d,\bm{0})}&\dots&\log\frac{\pi_{1K}(d,\bm{x})}{\pi_{1K}(d,\bm{0})}\\ \log\frac{\pi_{21}(d,\bm{x})}{\pi_{21}(d,\bm{0})}&\log\frac{\pi_{22}(d,\bm{x})}{\pi_{22}(d,\bm{0})}&\dots&\log\frac{\pi_{2K}(d,\bm{x})}{\pi_{2K}(d,\bm{0})}\\ \vdots&\vdots&\ddots&\vdots\\ \log\frac{\pi_{K1}(d,\bm{x})}{\pi_{K1}(d,\bm{0})}&\log\frac{\pi_{K2}(d,\bm{x})}{\pi_{K2}(d,\bm{0})}&\dots&\log\frac{\pi_{KK}(d,\bm{x})}{\pi_{KK}(d,\bm{0})}\end{bmatrix}.$
and the functions $\bm{\Gamma}\colon\{0,1\}\to\mathbb{R}^{n\times n}$ and $\bm{\Lambda}_{s}\colon\{0,1\}\to\mathbb{R}^{n\times n}$ $(s=1,\ldots,p)$ such that
@@
$\bm{\Delta}(\bm{x})\coloneqq\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}$
@@
$\Lambda_{ik}^{(s)}$ $\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)=$
$\left(\sum_{\bm{x}\in\mathcal{X}}\sum_{d\in\{0,1\}}\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\,\bm{\xi}^{(s)}\,\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\xi}^{(s)}\right)+\left(\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\right)^{\top}\,\bm{\xi}^{(s)}\,\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\xi}^{(s)}\right)^{\top}\right)_{ik}$
@@
$\bm{\pi}^{(s+1)}(1,\bm{x})=$$\big{\{}\left(\bm{\xi}^{(s+1)}\right)^{\top}\bm{g}\circ\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\xi}^{(s+1)}\right)\big{\}}\oslash$
$\left\{\left(\bm{\xi}^{(s+1)}\right)^{\top}\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\xi}^{(s+1)}\right)\right\},$
enabling the evaluation of $\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\xi}^{(s)}\right)$ for $\bm{x}\in\mathcal{X}$. Following Dahbura et al. (2021), we can still evaluate this matrix by sparse matrix operations.
In sum, we have shown how to compute the terms $\Omega_{ik}^{(s)}\left(\bm{0},\bm{0},\bm{\xi}^{(s)}\right)$ and $\Lambda_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ through matrix multiplications. Plugging in these terms into (A.1.2) enables a fast computation of $\Omega_{ik}^{(s)}(\bm{g},\bm{x},\bm{\xi}^{(s)})$, which is the computational bottleneck for evaluating the quadratic term needed to update $\bm{\xi}^{(s)}$ to $\bm{\xi}^{(s+1)}$.
A.2 STEP 2: Estimation of structural utility parameters
@@
$=$ $\arg\max_{\bm{\alpha},\bm{\beta},\gamma}\;\sum_{i\neq j}^{n}\,\left[g_{ij}\log p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}})+(1-g_{ij})\log(1-p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}}))\right]$
@@
Appendix B Additional figures and estimates
Figure A1: Value of the lower bound of the likelihood in the first step of our estimation-
Figure A2: Comparison of $k$-step systemicness for the 1% of repositories in the Rust Cargo ecosystem with the highest number of dependent repositories (in-degree). The $k$-step systemicness measures how many repositories the average vulnerable repository renders vulnerable $k$ steps away. The solid black line shows the average $k$-step systemicness when no node is protected, while the dashed and dot-dashed black lines show the average $k$-step systemicness when the 10 and 100 most systemic repositories are perfectly protected against vulnerabilities, respectively. The dashed and dot-dashed blue line shows the average $k$-step systemicness when systemicness for protected nodes are measued as the number of depdents a repository has (in-degree).

### v2

@@ Appendix B Exploiting sparse matrix operations for updating variational parameters
$\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)~{}$ $=~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$=\sum_{\bm{x}\in\mathcal{X}}\sum_{j\neq i}^{N}\sum_{l=1}^{K}g_{ij}\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)$
$+\,g_{ji}\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)+(1-g_{ij})\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\,\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)$
+ $+\,1-g_{ji})\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\,\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)$
@@
$\bm{\Pi}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\begin{bmatrix}\log\frac{\pi_{11}^{(s)}(d,\bm{x})}{\pi_{11}^{(s)}(d,\bm{0})}&\log\frac{\pi_{12}^{(s)}(d,\bm{x})}{\pi_{12}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{1K}^{(s)}(d,\bm{x})}{\pi_{1K}^{(s)}(d,\bm{0})}\\ \log\frac{\pi_{21}^{(s)}(d,\bm{x})}{\pi_{21}^{(s)}(d,\bm{0})}&\log\frac{\pi_{22}^{(s)}(d,\bm{x})}{\pi_{22}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{2K}^{(s)}(d,\bm{x})}{\pi_{2K}^{(s)}(d,\bm{0})}\\ \vdots&\vdots&\ddots&\vdots\\ \log\frac{\pi_{K1}^{(s)}(d,\bm{x})}{\pi_{K1}^{(s)}(d,\bm{0})}&\log\frac{\pi_{K2}^{(s)}(d,\bm{x})}{\pi_{K2}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{KK}^{(s)}(d,\bm{x})}{\pi_{KK}^{(s)}(d,\bm{0})}\end{bmatrix}$
and the functions $\bm{\Gamma}\colon\{0,1\}\to\mathbb{R}^{n\times n}$ and $\bm{\Lambda}_{s}\colon\{0,1\}\to\mathbb{R}^{n\times n}$ $(s=1,\ldots,p)$ defined as
@@
$\bm{\Delta}(\bm{x})\coloneqq\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}},$
@@
$\Lambda_{ik}$ $\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)=$ $\left(\sum_{\bm{x}\in\mathcal{X}}\sum_{d\in\{0,1\}}\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\,\bm{\Xi}^{(s)}\,\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)+\left(\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\right)^{\top}\,\bm{\Xi}^{(s)}\,\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)^{\top}\right)_{ik}.$
@@
$\bm{\pi}^{(s+1)}(1,\bm{x})=$ $\big{(}\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{g}\circ\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\Xi}^{(s+1)}\right)\big{)}\oslash$
$\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\Xi}^{(s+1)}\right)\right),$
enabling the evaluation of $\bm{\Pi}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)$ for $\bm{x}\in\mathcal{X}$. Following Dahbura et al. (2021), we can still evaluate this matrix by sparse matrix operations.
In sum, we have shown how to compute the terms $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ and $\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ through matrix multiplications. Plugging in these terms into (B) enables a fast computation of $\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})$, which is the computational bottleneck for evaluating the quadratic term needed to update $\bm{\Xi}^{(s)}$ to $\bm{\Xi}^{(s+1)}$.
B.1 STEP 2: Estimation of structural utility parameters
@@
$=$ $\arg\max_{\bm{\alpha},\bm{\beta},\gamma}\;\sum_{i\neq j}^{n}\,\left(g_{ij}\log p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}})+(1-g_{ij})\log(1-p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}}))\right)$
@@
Appendix C Additional figures and estimates
C.1 Comparison with other models
In Figure A1 we show the simulated statistics generated by a model with $K=1$. In aggregate, this model performs very poorly. This suggests that the unobserved heterogeneity can help fit the model better.
Figure A1: Model fit, ERGM ($K=1$)
We compare our model with $K=10$ unobserved types to a stochastic blockmodel with $K=10$ unobserved types. In essence this is the same as our model, but we exclude the externality. As expected the stochastic blockmodel does a worse job predicting the externality within blocks, as shown in Figure A2.
Figure A2: Model fit, stochastic blockmodel with $K=10$
We also did many robustness checks to choose the number of types $K$. In Figures A3 and A4 we report the boxplot for the fit of models with $K=5$ and $K=15$, respectively.
Figure A3: Model fit, model with $K=5$
Figure A4: Model fit, model with $K=15$
From the graph we can conclude that increasing or decreasing the number of types does not seem to help in estimating the externality.
Table A1: Parameter estimates and standard errors.
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-5.703$ $-7.979$
(0.274) (0.006)
$\log(n)\times$ Edges ($\alpha_{2})$ $-0.549$
Maturity ($\beta_{1}$) 1.767 $-0.222$
(0.303) (0.011)
$\log(n)\times$ Maturity ($\beta_{1}$) $-0.12$
Popularity ($\beta_{1}$) 1.071 $-0.07$
(0.308) (0.011)
$\log(n)\times$Popularity ($\beta_{1}$) $-0.117$
Size ($\beta_{1}$) 1.92 0.118
(0.303) (0.01)
$\log(n)\times$Size ($\beta_{1}$) $-0.178$
Table A2: Parameter estimates and standard errors ($K=5$).
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-4.893$ $-7.45$
(0.263) (0.006)
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.569$
Externality ($\gamma$) $-0.258$
Maturity ($\beta_{1}$) 0.711 $-0.267$
(0.302) (0.012)
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.012$
Popularity ($\beta_{1}$) 1.316 0.104
(0.31) (0.011)
$\log(n)\times$Popularity ($\beta_{1}$) $-0.111$
Size ($\beta_{1}$) 1.324 $-0.073$
(0.307) (0.011)
$\log(n)\times$Size $\left(\beta_{1}^{\ell}\right)$ $-0.142$
Table A3: Parameter estimates and standard errors ($K=15$).
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-4.181$ $-8.28$
(0.305) (0.006)
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.765$
Externality ($\gamma$) 0.135
Maturity ($\beta_{1}$) 3.201 $-0.198$
(0.321) (0.011)
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.307$
Popularity ($\beta_{1}$) 1.58 0.117
(0.287) (0.01)
$\log(n)\times$Popularity $\left(\beta_{1}^{\ell}\right)$ $-0.15$
Size ($\beta_{1}$) 0.369 $-0.074$
(0.29) (0.01)
$\log(n)\times$Size $\left(\beta_{1}^{\ell}\right)$ $-0.048$
ERGM Model
Estimate
(Std. Error))
Edges $-7.961$
$(0.007)$
Externality $-0.178$
$(0.002)$
Maturity $-0.131$
$(0.010)$
Size $0.127$
$(0.010)$
Popularity $-0.066$
$(0.010)$
Table A4: Parameter estimates and standard errors (K=1)